A person drops a cylindrical steel bar ( Y = 1.600 × 10 11 Pa ) from a height of 4.00 m (distance between the floor and the bottom of the vertically oriented bar). The bar, of length L = 0.940 m , radius R = 0.00550 m , and mass m = 1.900 kg , hits the floor and bounces up, maintaining its vertical orientation. Assuming the collision with the floor is elastic, and that no rotation occurs, what is the maximum compression of the bar?

Maximum compression: ____m
Answer #1

__________________________________________


Y = 1.6x10^11 Pa
h = 4 m
L = 0.94 m
R = 0.0055 m
m = 1.9 kg

maximum compression of the bar = dx
then from conservation of energy

mgh = U
also
U = P^2L/2YA
but
dx = PL/YA
hence
U = dx^2*YA/2L

hence
mgh = dx^2*Y*pi*R^2/2L
1.9*9.81*4 = dx^2*1.6x10^11*pi*0.0055^2/2*0.94
dx = 0.00303615 m ( maximum deflection)

__________________________________________

Please rate
if any mistake in this answer please comment i will clarify your doubt . thank you